Shift report 20011016 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 16 2001 - 08:26:50 EDT

  • Next message: Flemming Videbaek: "Fw: BRAHMS NIM Paper "final" revision"

    ------------------------------------------------------------
    Date: Tue Oct 16 08:26:49 2001, Shift: 00:00-08:00
    Supervisor: Y. K. Lee, with: 
    Log book#10 pages 15 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    1. Ramping was difficult. There was only one store of
       three and 1/2 hours and the beam was not satisfactorily
       steered to BRAHMS. The ZDC rate was 220 /sec.
    
    2. MRS @45 deg, 4B; FS @4/4 deg, 1/5B.
    	Run 5367 had 225K trigger-1.
            Run 5368 had 91K trigger-1.
       
       Pulser and source runs
            Run 5365 is Detector pedestal run.
            Run 5366 is TPC pedestal run.
            Run 5369 is Drift Velocity Monitor run.
    
    3. Detector problems:
    	ZDC still has the Right Module 1 Low missing.
    	BB has ADC Left Module 37 missing.
    	BB Small Right ADC 25 has small signal (1/5).
    	TOFW TOP 92 slat missing.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 16 2001 - 08:27:33 EDT